### v2

# Disentangled Latent Energy-Based Style Translation: An Image-Level Structural MRI Harmonization Framework

Abstract
Brain magnetic resonance imaging (MRI) has been extensively employed across clinical and research fields, but often exhibits sensitivity to site effects arising from non-biological variations such as differences in field strength and scanner vendors. Numerous retrospective MRI harmonization techniques have demonstrated encouraging outcomes in reducing the site effects at image level. However, existing methods generally suffer from high computational requirements and limited generalizability, restricting their applicability to unseen MRIs. In this paper, we design a novel disentangled latent energy-based style translation (DLEST) framework for unpaired image-level MRI harmonization, consisting of (a) site-invariant image generation (SIG), (b) site-specific style translation (SST), and (c) site-specific MRI synthesis (SMS). Specifically, the SIG employs a latent autoencoder to encode MRIs into a low-dimensional latent space and reconstruct MRIs from latent codes. The SST utilizes an energy-based model to comprehend global latent distribution of a target domain and translate source latent codes toward the target domain, while SMS enables MRI synthesis with a target-specific style. By disentangling image generation and style translation in latent space, the DLEST can achieve efficient style translation. Our model was trained on T1-weighted MRIs from a public dataset (with 3,984 subjects across 58 acquisition sites/settings) and validated on an independent dataset (with 9 traveling subjects scanned in 11 sites/settings) in four tasks: histogram and feature visualization, site classification, brain tissue segmentation, and site-specific structural MRI synthesis. Qualitative and quantitative results demonstrate the superiority of our method over several state-of-the-arts.
keywords:
MRI harmonization , style translation , MRI synthesis , energy-based model
journal: arXiv
Figure 1: The proposed disentangled latent energy-based style translation (DLEST) framework for MRI harmonization at the image level, consisting of (a) site-invariant image generation that encodes images into low-dimensional latent space and reconstruct images based on latent codes, (b) site-specific style translation that facilitates implicit style translation within the latent space, and (c) site-specific MRI synthesis that generates diverse synthetic MRIs with a given target site style.
1 Introduction
Structural magnetic resonance imaging (MRI) has found extensive application in both clinical diagnostics and research studies, primarily due to its ability to offer high spatial resolution in a non-invasive manner [10, 1, 41]. Large-scale data initiatives and neuroimaging studies often utilize brain MRI data pooled from various acquisition sites/settings to enhance subject cohort and boost the statistical power of learning models [2, 40, 33]. However, models directly trained on these multi-site datasets often encounter difficulties in achieving convergence [12, 38] and in generalizing to new data, due to the sensitivity of MRI to non-biological variations, known as site effects. The site effect can be attributed to a variety of factors, including differences in field strengths, scanner vendors, scanning protocols, and hardware upgrade and maintenance. [12, 42, 16]. While multi-site studies, such as Alzheimer’s Disease Neuroimaging Initiative [17], aim to mitigate site effects during data acquisition through unified scanning protocols, certain variations are inevitable, like those caused by scanner maintenance or software upgrades. Additionally, many studies rely on the retrospective pooling of multi-site data, when there was no coordination among different sites during image acquisition [9]. One possible solution to address such a challenge is to retrospectively harmonize the imaging data as a preprocessing step [3, 49].
Existing methods for MRI harmonization can generally be classified into two categories: feature-level approaches and image-level approaches. Feature-level techniques, such as ComBat [18] and ComBat-GAM [30], typically utilize empirical Bayes models and pre-extracted biological or radiomic features from a set of pre-defined regions-of-interest (ROIs). Nonetheless, the effectiveness of these methods is largely contingent on the size of datasets and the quality of feature extraction, which restricts their versatility across applications that employ diverse MRI features [3, 2]. On the other hand, image-level harmonization is designed for a wider range of downstream applications. It is not reliant on pre-defined MRI features, which is the primary objective of this paper.
For image-level harmonization, many deep generative models have been employed in recent studies [48, 5, 23, 49, 3]. These methods, rather than harmonizing a set of pre-extracted MRI features, approach the harmonization problem as a pixel-to-pixel translation task at the image level. The objective is to alter the style of a source MRI to match a target MRI while preserving the original content. In this context, the style information refers to high-level features like intensity difference, contrast variations, and signal-to-noise ratio, while content information pertains to low-level image features such as anatomical structures, contours, and edges. Several state-of-the-art (SOTA) methods utilize generative adversarial networks (GANs) for image-level MRI harmonization. For instance, CycleGAN [48] imposes cycle-consistency constraints in its loss function to enforce content preservation during unpaired style translation. StarGAN [5] employs separate encoders to learn site-specific style codes and injects them into images via adaptive instance normalization. ImUnity [3] incorporates additional encoder networks as extra constraints for learning latent representations to remove the site bias and retain clinical information. In addition to GAN-based models, some recent works utilize encoder-decoder models to learn disentangled latent representations for anatomical structure and contrast to perform image-level MRI harmonization [50, 7]. However, existing image-level methods usually utilize deep networks to separately learn style and content encoding in image space, leading to significantly increased training time and computational cost due to the optimization of numerous network parameters [14]. Furthermore, many approaches need to retrain their models entirely when applied to unseen MRI data, therefore limiting their practical generalizability [3].
In this paper, we propose a novel disentangled latent energy-based style translation (DLEST) framework for unpaired structural MRI harmonization at the image level. As illustrated in Fig. 1, the DLEST includes (a) a site-invariant image generation module, using a latent autoencoder for encoding and generating MRIs from a lower-dimensional latent space, (b) a site-specific style translation module, using an energy-based model for implicit style translation through latent space, and (c) a site-specific MRI synthesis module for generating synthetic MRIs with given site-specific styles. Through disentangling image generation and style translation in latent space, the DLEST can achieve efficient and generalizable MRI harmonization. It can also be extended to synthesize MRIs with site-specific styles and diverse anatomical details, without extra model training.
An initial version of our work has been introduced in MLMI 2023 [45]. This journal paper introduces several advancements, including 1) extension of our framework for site-specific MRI synthesis, with qualitative and quantitative evaluations, 2) a thorough literature review on image-level harmonization methods and the application of energy-based models for generation tasks, 3) adding comparison with a recent state-of-the-art method, 4) experimentation with specific loss functions for content and style regularization, 5) adding comparative visualization of brain tissue segmentation results, 6) including experimental analysis on the influence of target site selection, and 7) adding quantitative comparison of computational cost across all learning methods.
The key contributions of this work are outlined as follows.
  • 1.
    We propose a novel disentangled latent energy-based style translation framework for unpaired MRI harmonization at the image level, achieving more computationally efficient style translation in low-dimensional latent space.
  • 2.
    Our framework can generalize to new data without model retraining since the image generation module can directly encode and reconstruct MRIs without site specification, while the style translation module can be fine-tuned during the testing stage with minimal computational cost. Once trained for harmonization, our method can synthesize MRIs with the style of a specified site.
  • 3.
    We incorporate an energy-based model to capture the latent data distribution of target MRI scans for harmonization. This enables our model to more effectively account for intra-site variations within the target domain and avoids pre-defining a reference image for harmonization.
  • 4.
    Our framework is rigorously evaluated on two public datasets with T1-weighted MRIs across four tasks, with qualitative and quantitative results underscoring the superiority of our DLEST over SOTA methods.
2 Related Work
2.1 Image-Level MRI Harmonization
Inspired by the image-to-image translation task in the natural imaging domain, recent studies have employed the generative adversarial network (GAN) models to tackle medical data harmonization problems on the image level [48, 5, 23]. These methods engage the generator and discriminator networks in an adversarial game, where the generator creates synthetic images resembling the real dataset distribution, and the discriminator differentiates between synthetic and real images [48]. For instance, CycleGAN introduces a cycle-consistency constraint in its loss function for unpaired image translation and content (anatomical structure) preservation [48]. Style-encoding GAN [23], inspired by StarGAN-V2 [5], further separates the content and style encoding in the latent space, allowing the site-specific style code to be learned and injected into images via adaptive instance normalization. ImUnity [3] modifies the GAN structure by adding a site/scanner unlearning module to encourage the encoder to learn domain-invariant latent representations. These advancements have contributed to the continual improvement of GAN-based harmonization methods.
In addition to GAN-based models, recent studies have introduced an alternative approach that employs encoder-decoder networks to disentangle anatomical and contrast information in latent space. For instance, CALAMITI [49] first uses T1- and T2-weighted (T1/T2-w) MRI pairs to learn global latent codes containing anatomical and contrast information, and then disentangles style and content latent codes via separate encoders and decoders. Similarly, Dewey et al.  [7] leverage T1-w and T2-w image pairs to attain a disentangled latent space, comprising high-resolution anatomical and low-dimensional contrast components via a Randomization block. This block allows generating MRIs with identical anatomical structures but varying contrast. Zuo et al.  [50] enhance this approach by ablating the need for paired MRI sequences. Instead, they employ 2D slices from axial and coronal views of the same MRI to provide the same contrast but different anatomical information.
These existing methods usually involve simultaneous training of multiple encoders, decoders, or sub-networks for separated style and content encoding, resulting in increased computational and time costs for model training. Besides, some methods have limited generalizability on new MRI as their models need to be retrained entirely on new datasets. Moreover, certain existing studies necessitate paired MRI scans for model training, like traveling subjects or multiple MRI sequences, which are often unavailable in practice. In this work, we aim to enhance the efficiency of image-level harmonization and the model’s generalizability to new data, while also expanding its capacity for site-specific image synthesis.
2.2 Energy-Based Models for Image Generation
Generative models are a category of probabilistic models that represent the joint probability distribution of both observed and target data. For the image generation task, a model observes image samples or their latent representations and learns to capture the true data distribution to generate new images that resemble the observed samples [27, 8]. Probabilistic models employ statistical inference methods, such as maximum likelihood estimation (MLE), to estimate parameters that maximize the likelihood of the observed data under an assumed distribution [26]. However, these models need to be correctly normalized. This process can sometimes involve the evaluation of intractable integrals over all possible variable configurations, which can be computationally challenging. Therefore, LeCun et al.[21] use energy-based learning as an alternative framework for both discriminative and generative tasks. Energy-based models (EBMs) capture data dependencies by associating a scalar energy value (a measure of compatibility) to each configuration of variables [21]. Unlike probabilistic models, EBMs do not require normalization. For generative tasks, the training phase of EBMs involves learning an appropriate energy function that assigns low energy values to observed data within the real data distribution and high energy values otherwise, during which data is typically sampled via Markov Chain Monte Carlo (MCMC) sampling methods. During inference, the EBMs generate data by exploring the energy surface to sample configurations that minimize the energy function.
Over the years, many studies have employed EBMs for image generation and improved on various aspects. Du et al.[8] propose to utilize Langevin Dynamics, a gradient-based MCMC method, for more efficient sampling and mixing on high-dimensional image domains. Pang et al.[27] further improves the sampling efficiency and the MCMC mixing by combining an EBM with a top-down network (decoder) in the low-dimensional latent space. The EBM serves to refine the simple isotropic Gaussian prior to be closer to the data distribution, leading to improved quality during image generation. Xiao et al.[46] propose a symbiosis of Variational Autoencoder (VAE) and EBM to leverage the advantages of both. The VAE captures the overall model structure of the data distribution and provides an improved latent initialization of the MCMC sampling, which is used in EBM. The EBM, on the other hand, can exclude non-data-like samples and refine the latent representation for VAE to decode. A recent work [47] utilizes this symbiotic combination of EBM and autoencoder for the task of male-to-female face translation. In this context, the EBM learns the latent distribution of a target image domain and translates the source latent representation towards this target distribution. The underlying assumption is that two image domains share certain common latent representations. The EBM is then able to implicitly translate latent codes that encapsulate domain differences while preserving shared latent codes that signify domain similarities. Inspired by the success of EBMs in natural image-to-image translation, our method adopts a similar assumption. We posit that MRIs from different domains share some common latent representations that convey content (anatomical) information. We propose training an EBM to translate the site-specific style information from one domain to another within a low-dimensional latent space.
3 Materials and Methodology
3.1 Materials and Image Preprocessing
3.1.1 Datasets
Our study utilizes two public datasets, including (1) Open Big Healthy Brains (OpenBHB) [9] and (2) Strategic Research Program for Brain Science (SRPBS) [39]. Specifically, the OpenBHB, a large-scale brain MRI dataset, comprises 3,984 T1-weighted MRIs from healthy subjects acquired from over 58 centers globally. The SRPBS dataset contains 108 T1-weighted structural MRIs acquired from 9 healthy traveling subjects across 12 scanners/settings. In this work, MR scans of 9 subjects from 11 acquisition sites/settings in SRPBS are used, while one duplicated site (YC2) is removed. Further details can be found online [^1] https://baobablab.github.io/bhb/, [^2] https://www.synapse.org/#!Synapse:syn22317076/wiki/605026.
3.1.2 Data Preprocessing
For these two datasets, each MRI scan undergoes minimal preprocessing using FSL ANAT pipeline [36], which encompasses bias field correction, brain extraction, and 9 degrees of freedom linear registration to the MNI-152 template with $1\,mm$ isotropic resolution. Brain tissue segmentation is performed on SRPBS images using FSL FAST pipeline, with each brain segmented into white matter (WM), gray matter (GM), and cerebrospinal fluid (CSF). For each MRI from OpenBHB, we select 10 slices in axial view from the center of each MRI. We utilize the original training and test data split specified in [9], which eventually results in $32,270$ 2D MRI slices for training and $7,570$ MRI slices for validation. For SRPBS with limited traveling subject MRIs, we choose 14 slices from each volume, resulting in $1,386$ 2D MRI slices. Additionally, all these 2D slices are zero-padded to have the size of $256\times 256$.
3.2 Proposed Methodology
3.2.1 Problem Formulation
In this work, we reformulate image-level MRI harmonization as a domain-level translation task, which is different from previous studies that treat each MRI scan as a unique domain and concentrate on individual source-to-reference image translation [23, 3]. In the context of multi-site MRI studies, we consider a single site or scanning setting for MRI acquisition as a unique domain. This is justified by numerous studies demonstrating that MRIs scanned in the same site under the same setting can exhibit intra-site variations that result in different styles within one domain [6, 37]. Consequently, image-level translation using only one reference image may not account for these variations, failing to capture true data distributions of a site/domain. In this work, we propose to employ all MRIs from a target site as reference images, aiming to better encapsulate the stylistic features of target data. Overall, we aim to translate MRIs within a source domain $X$ to have the imaging style (i.e., non-biological variation) of MRIs in a target domain $Y$.
[…]
3.2.2 Site-Invariant Image Generation (SIG)
As shown in Fig. 1 (a), the SIG is designed to encode MRI scans into a lower-dimensional latent space and subsequently decode reconstructed MRIs from these latent codes. Importantly, this process is trained on the OpenBHB training set, independently of any site specifications, and thus, does not require any site labels or paired images. Specifically, the SIG module takes as input a real MRI ${\bf x}\sim\mathbb{P}_{Data}$ and a random noisy image $z$ drawn from a Gaussian distribution $\mathcal{N}(0,1)$, where $\mathbb{P}_{Data}$ denotes the real data distribution. A network $\bm{F}$ learns a deterministic mapping to convert $z$ into a latent code [29], represented as $\mathcal{W}={F}(z)$. Following this, we further design a stochastic decoding network $\bm{G}$ to generate a synthetic MR image based on the latent code $\mathcal{W}$ and an independent noise $\eta\sim\mathcal{N}(0,1)$, represented as $\tilde{{\bf x}}={G}(\mathcal{W},\eta)$. A network $\bm{E}$ is then trained to encode the synthetic image $\tilde{{\bf x}}$ back into the latent space, resulting in a reconstructed latent code $\tilde{\mathcal{W}}$. This entire process can be symbolized as $\tilde{\mathcal{W}}=E(\tilde{{\bf x}})=E\circ G\circ F(z)$, where the symbol $\circ$ denotes function composition. The encoder $\bm{E}$ also can transform a real image ${\bf x}$ into its corresponding latent code, denoted as $\mathcal{R}=E({\bf x})$. Lastly, a latent discriminator, denoted as $\bm{D}$, receives a latent code and produces a probability value that signifies whether the latent code is derived from a synthetic or real MRI. The networks $\bm{F}$, $\bm{G}$, $\bm{E}$, and $\bm{D}$ are trained using an adversarial approach, enabling the SIG module to generate latent codes that closely resemble those derived from real MRIs. The latent code $\mathcal{R}$, derived from the real image, is also decoded back to a reconstructed image using the decoder $\bm{G}$, aiming to preserve the content (e.g., anatomical structures) of the real MRI.
To uphold the symmetry of the encoder and decoder in latent space, we aim to ensure that $\mathcal{W}=\tilde{\mathcal{W}}$, accomplished by minimizing the following latent autoencoder error loss:
$\mathcal{L}_{lae}^{E,G}=\Delta_{1}(F\lVert E\circ G\circ F)=\mathbb{E}_{z\sim\mathbb{P}_{z}}[\lVert F(z)-E\circ G\circ F(z)\rVert_{2}^{2}]$ (1)
Through this process, the network pair $\bm{E}$ and $\bm{G}$ can be interpreted as a latent autoencoder that encodes and reconstructs the latent code $\mathcal{W}$. This differs from conventional autoencoders, which typically operate in high-dimensional image space [20]. The generation of images via autoencoding low-dimensional latent codes enhances computational efficiency.
To maintain the anatomical structure preservation of the latent autoencoder $\bm{E}$ and $\bm{G}$, a pixel-wise content loss is introduced to compute the difference between the original real image ${\bf x}$ and its reconstructed counterpart $\tilde{{\bf x}}$, as follows:
$\mathcal{L}_{pix}^{E,G}=\Delta_{2}({\bf x}\lVert\tilde{{\bf x}})=\mathbb{E}_{x\sim\mathbb{P}_{Data}}[\lVert{\bf x}-G\circ E({\bf x})\rVert_{1}]$ (2)
which motivates the model to minimize content change during MRI reconstruction. To generate faithful synthetic MRIs, we propose to jointly optimize $\bm{F}$, $\bm{G}$, $\bm{E}$, and $\bm{D}$ through an adversarial loss, which is formulated as:
$\mathcal{L}_{adv}^{E,D}=\Phi(D\circ E\circ G\circ F(z))+\Phi(-D\circ E({\bf x}))+R_{reg}$ (3)
$\mathcal{L}_{adv}^{F,G}=\Phi(-D\circ E\circ G\circ F(z))$ (4)
where $\Phi$ is the $softplus$ function defined as $\Phi(t)=\log(1+\exp(t))$, $R_{reg}$ is a zero-centered gradient penalty term defined as $\frac{\gamma}{2}\mathbb{E}_{x\sim\mathbb{P}_{Data}}[\lVert\nabla D\circ E({\bf x})\rVert^{2}]$ [29].
Overall, the final objective functions of SIG are as follows:
${\operatorname{min}_{\{F,G\}}}\,{\operatorname{max}_{\{E,D\}}}\,\mathcal{L}_{adv}^{E,D}+\mathcal{L}_{adv}^{F,G},~{}~{}{\operatorname{min}_{\{E,G\}}}\,\mathcal{L}_{lae}^{E,G}+\mathcal{L}_{pix}^{E,G}$ (5)
3.2.3 Site-Specific Style Translation (SST)
As depicted in Fig. 1 (b), the proposed SST consists of (1) an autoencoder that includes the networks $\bm{E}$ and $\bm{G}$ trained in the SIG module, and (2) an energy-based model (EBM) that performs latent-space style translation. Given an MRI ${\bf x}$ from source site $X$ and an MRI ${\bf y}$ from target site $Y$, the encoder $\bm{E}$ initially transforms them into their respective latent codes, denoted as $Z_{x}=E({\bf x})$ and $Z_{y}=E({\bf y})$, respectively. These latent codes are then input into the EBM, which produces the translated latent code $Z_{x\to y}$. This translated latent code is subsequently decoded by $\bm{G}$ to produce the translated MRI. The EBM is also utilized to reverse-translate $Z_{x\to y}$, resulting in a recovered source latent code, denoted as $\tilde{Z}_{x}$. This reconstructed latent code is compared with the original $Z_{x}$ to ensure the latent cycle-consistency, which prevents mode collapse during the unpaired style translation process. The specifics of the EBM are discussed as follows.
[…]
$\begin{split}\nabla_{\theta}\mathcal{L}_{EBM}&=-\frac{\partial}{\partial\theta}L(\theta)\\ &=\mathbb{E}_{Z_{y}\sim\mathbb{P}_{Data}}\left[\frac{\partial}{\partial\theta}\mathcal{E}_{\theta}(Z_{y})\right]-\mathbb{E}_{\tilde{Z}_{y}\sim\mathbb{P}_{\theta}}\left[\frac{\partial}{\partial\theta}\mathcal{E}_{\theta}(\tilde{Z}_{y})\right]\end{split}$ (6)
where the first expectation term $\mathbb{E}_{Z_{y}}$ corresponds to samples $Z_{y}$, which are drawn from the actual target data distribution $\mathbb{P}_{Data}(Y)$. On the other hand, the second expectation term $\mathbb{E}_{\tilde{Z}_{y}}$ corresponds to samples $\tilde{Z}_{y}$, which are drawn from the modeled distribution $\mathbb{P}_{\theta}$. This requires direct sampling from $\mathbb{P}_{\theta}$, an intractable model distribution. This can be approximated using a Markov chain Monte Carlo (MCMC) sampling method, namely the Stochastic Gradient Langevin Dynamics (SGLD) [44], an iterative method defined as follows:
$\tilde{Z}_{y}^{t+1}=\tilde{Z}_{y}^{t}-\frac{\eta^{t}}{2}\frac{\partial}{\partial\tilde{Z}_{y}^{t}}\mathcal{E}_{\theta}(\tilde{Z}_{y}^{t})+\sqrt{\eta^{t}}\epsilon^{t},\ \epsilon^{t}\sim\mathcal{N}(0,I)$ (7)
where $t=[0,\cdots,T]$ represents the sampling steps with a step size $\eta$ and $\epsilon$ is a Gaussian noise that is injected to account for data uncertainty and to improve sample convergence [21]. The initial $\tilde{Z}_{y}^{0}$ are latent codes sampled from the source site $X$, denoted as $\tilde{Z}_{y}^{0}=Z_{x}=E({\bf x})$. The SGLD sampling process involves iteratively updating the initial source latent codes using negative gradients of the energy function $\mathcal{E}_{\theta}$. This sampling process can be interpreted as evolving original source latent codes towards a configuration with lower energy, thereby aligning them more closely with the true target latent distribution. In contrast to source-to-reference image translation [23, 48, 5], EBM effectively traverses the latent space of the entire target domain, allowing a comprehensive consideration for intra-site variations within the target domain. After $T$ sampling steps (via SGLD), we can decode $\tilde{Z}_{y}^{T}$ through $\bm{G}$ to generate a translated MRI, denoted as $\tilde{{\bf y}}=G(Z_{x\to y})=G(\tilde{Z}_{y}^{T})$.
Considering that the EBM does not differentiate the content code and the style code explicitly during style translation, we further introduce a latent content loss into the EBM for regularization measure. This approach is designed to prompt the model to minimally alter the latent code during style translation, thereby preserving the overall content. This loss function measures the difference between the translated latent code $Z_{x\to y}$ and the original source latent code $Z_{x}$, which is formulated as:
$\mathcal{L}_{con}=\mathbb{E}_{Z_{x}\sim\mathbb{P}_{Data},Z_{x\to y}\sim\mathbb{P}_{\theta}}[\lVert Z_{x}-Z_{x\to y}\rVert_{1}]$ (8)
Instead of using the latent content loss in Eq. (8), one can also explicitly compute the content loss and the style loss during the latent style translation, which will be discussed in Section 5.3.
To guarantee the learned translation is bijective and to further avoid mode collapse during unpaired latent style translation, we introduce a latent cycle-consistency loss, defined as:
$\mathcal{L}_{cyc}=\mathbb{E}_{Z_{x}\sim\mathbb{P}_{Data}}[\lVert Z_{x}-\mathbb{P}_{\theta}^{-1}\circ E\circ G\circ\mathbb{P}_{\theta}(Z_{x})\rVert_{1}]$ (9)
where $\mathbb{P}_{\theta}^{-1}$ represents the same EBM as $\mathbb{P}_{\theta}$, but it operates with an inverse SGLD sampling process. This means that it will update the latent code using positive gradients of the energy function, in contrast to the negative gradients used by $\mathbb{P}_{\theta}$.
With the integration of Eqs. (6), (8)-(9), the full objective function of SST can be written as:
$\mathcal{L}(\theta)=\mathcal{L}_{EBM}+\alpha\mathcal{L}_{con}+\beta\mathcal{L}_{cyc}$ (10)
where $\alpha$ and $\beta$ serve as hyperparameters that determine the relative contributions of the three terms in the equation. The SST can be trained or fine-tuned for any target style translation with a small sample size for each target site (i.e., $\sim$100100100100 MRI slices in our experiments). The fine-tuning process on a new target site proves to be relatively efficient in terms of time and computation resources. In our experiments, it takes approximately half an hour and utilizes 10G (out of 24G) of GPU memory on an RTX 3090 GPU. More details are discussed in Section 5.5. By disentangling image generation via SIG and style translation via SST in the low-dimensional latent space, it is expected to achieve efficient and generalizable MRI harmonization.
3.2.4 Site-Specific MRI Synthesis (SMS)
In the proposed SST module, we can substitute the original source MRI and encoder $\bm{E}$ with a random Gaussian noise $z$ and the mapping network $\bm{F}$ to perform site-specific MRI synthesis, as shown in Fig. 1 (c). In this case, as the network $\bm{F}$ has already been trained in the SIG to generate latent codes resembling brain MRIs, no further fine-tuning is required. The EBM will then translate these randomly generated latent codes into the target style for image synthesis. Unlike many MRI synthesis models that only generate MRI with random anatomical structures and styles, our framework can synthesize MRIs with a consistent intensity distribution that matches a given image domain, this can be effectively used as a data augmentation method to pre-train models that require large number of MRI samples without introducing non-biological variations caused by difference in imaging style.
Figure 2: Histogram comparison of 10 source sites and a target site (COI) across all 9 traveling subjects from the SRPBS dataset. The first plot shows pre-harmonization histograms, while the subsequent plots depict post-harmonization histograms by each competing method and our method DLEST, respectively. WM: white matter; GM: gray matter; CSF: cerebrospinal fluid.
3.2.5 Implementation
The implementation of SIG incorporates similar settings as outlined in [29], with structural modifications made to accommodate our specific datasets and tasks. Specifically, both the mapping function $\bm{F}$ and the latent discriminator $\bm{D}$ are implemented using multi-layer perceptrons (MLPs). The former consists of eight layers, while the latter is composed of three layers. Two networks $\bm{E}$ and $\bm{G}$ in the latent autoencoder are composed of seven encoding/decoding blocks. Each of these blocks incorporates an increasing number of filters, ranging from 64 to 256 filters. All latent codes have a dimension of 512 in the experiments. We train the SIG module using an Adam [19] optimizer in an unsupervised training scheme. Compared to GANs [48, 5], SIG’s training cost is lower due to two factors: (1) most of its losses are evaluated in low-dimensional latent space, and (2) no additional cost for cycle-consistency constraints in image space. Following the SIG training, we can directly use the latent autoencoder in SST training, during which $\bm{E}$ and $\bm{G}$ are not updated.
The training of the SST is efficient in both training time and computational cost due to four factors. (1) The autoencoder $\bm{E}$ and $\bm{G}$ do not require additional training, even when deployed on previously unseen data. (2) The style translation process, facilitated by EBM, operates in a low-dimensional latent space with a dimension of 512, which is considerably smaller than the image space. (3) The energy function $\mathcal{E}_{\theta}$ in EBM is designed to be light-weight, also implemented as an MLP with 2 layers, each having a dimension of 2,048. (4) The SST can be fine-tuned with a small sample size at low computational cost, which will be discussed in Sections 5.4 and  5.5. In this way, our DLEST can achieve efficient imaging style translation and generalizable MRI generation through the disentanglement scheme.
4 Experiments
4.1 Experimental Settings
4.1.1 Competing Methods
We first compare DLEST with two non-learning methods, including (1) histogram matching (HM[34] with multiple experiments with randomly selected reference MRI from the target dataset, and (2) Spectrum Swapping-based Image-level Harmonization (SSIMH), which replaces certain low-frequency region of the source MRI with the corresponding low-frequency spectrum of the reference MRI, with a default swapping threshold 3. We use the DomainATM toolbox [13] to carry out two non-learning methods. We also compare DLEST against the following state-of-the-art (SOTA) methods for unpaired image-level harmonization.
(1) CycleGAN [48] is initially introduced for natural image-to-image translation. It enhances the capabilities of traditional GANs by integrating an important cycle-consistency constraint. This additional constraint enables the unpaired mapping from a source domain to a target domain. The CycleGAN has been utilized in several MRI harmonization studies [25, 4] and has become the foundational structure of many subsequent image-level harmonization methods.
(2) StarGAN [5] is initially introduced for multi-domain natural image synthesis task. Then, it was employed for MRI harmonization [23]. Building upon CycleGAN, the StarGAN method learns a separate latent style code for each target image and injects this code into the source image via adaptive instance normalization.
(3) ImUnity [3] is designed specifically for MRI harmonization. It employs a self-supervised variational autoencoder model, coupled with a confusion module for site/scanner-bias unlearning. In addition, it includes an optional biological preservation module to preserve clinical information, such as disease labels or diagnostic scores. In this work, we use the default form of ImUnity without this optional module, since the two datasets only consist of healthy control subjects.
In all three deep learning methods mentioned above, we typically adhere to the default setting of each competing method and make a dedicated effort to ensure other training hyperparameters align closely with the proposed DLEST framework. In DLEST, the two hyperparameters $\alpha$ and $\beta$ in Eq. 10 are empirically set as 1 and 100, respectively. For a fair comparison, we standardize the input size to be $256\times 256$ and set the latent dimension to 512, mirroring the specifications of the proposed DLEST. To train deep learning models (i.e., CycleGAN, StarGAN, ImUnity, and our DLEST), we use the OpenBHB dataset, due to its relatively larger sample size and abundant sites/settings. Following the data split in [9], $32,270$ training slices and $7,570$ validation slices are used, with manual inspection and quality control.
4.1.2 Evaluation Tasks
Four tasks are performed, including (1) histogram comparison and feature visualization, (2) site classification, (3) brain tissue segmentation, and (4) site-specific MRI synthesis. The SRPBS is used in Tasks 1, 3, and 4 because traveling subjects are available. The OpenBHB is used in Task 2 and Task 4.
4.2 Experimental Results
4.2.1 Task 1: Histogram Comparison & Feature Visualization
We first qualitatively evaluate our DLEST by comparing histograms and extracted features of MRIs before harmonization (Baseline) and post-harmonization produced by each competing method in Fig. 2. The validation is performed on SRPBS since each subject has ground-truth scans on multiple acquisition sites. To select a target site, we compare intra-site variations across each site in the SRPBS dataset by computing the mean peak signal-to-noise ratio (PSNR) for all image pairs within a given site. Given that the SRPBS encompasses all traveling subjects, each site shares an identical subject cohort, thereby containing consistent anatomical information. Consequently, a site yielding a higher intra-site PSNR score suggests a greater similarity in image style within that particular site. We treat COI as the target domain due to its low intra-site variation and harmonize the remaining 10 sites/settings to the COI. The influence of target site selection is elaborated in Section 5.4.
As observed in Fig. 2, before MRI harmonization (baseline), each site initially exhibits unique intensity distributions. In particular, three prominent intensity peaks, WM, GM, and CSF, are not aligned with the target domain. After harmonization, our DLEST framework demonstrates superior performance in aligning the histogram of the ten source sites with the target COI, particularly excelling in aligning the three main peaks. However, the three SOTA deep learning methods (i.e., CycleGAN, StarGAN, and ImUnity) cannot well align source peaks with the target ones, despite showing good peak alignment across the 10 source sites. Our success may be attributed to the model’s ability to learn the latent data distribution of the entire target domain via EBM, while the three deep learning methods do not effectively account for intra-site variation within the target domain. On the other hand, the two non-learning methods cannot produce good performance. For example, the HM method, despite achieving reasonable peak alignment with the target, exhibits a non-smooth artifact in the post-harmonization histogram. This could be due to the significant intensity disparity between the source MRIs and the target/reference MRI [22], caused by potential site effects.
For feature visualization, we use a pre-trained ResNet18 [15] model to extract features of the original MRIs and harmonized ones (via DLEST) across all sites and subjects. We then apply principal component analysis (with 100 principal components) to reduce the dimensionality of the extracted image features, and then utilize the t-SNE [24] to reduce the dimensionality of these features further and project them into a 2D space. Finally, we color-code each site with a unique color. The visualization results in Fig. 3 show that the extracted image features are initially clustered into 11 distinct groups. After harmonization by DLEST, these clusters become less distinguishable, implying that our method effectively eliminates site-related variations.
Figure 3: Visualization of MRI features of 11 sites across 9 subjects from SRPBS, with MRI (a) before and (b) after harmonization by DLEST. Each color denotes a specific site, while each point denotes extracted features from an MRI slice of a specific subject.
4.2.2 Task 2: Site Classification
In this task, we conduct a quantitative evaluation to determine the effectiveness of a harmonization method in eliminating site-related, non-biological variations. We begin by selecting a site (ID: 17) from OpenBHB with minimal intra-site variation as our target site. We then harmonize MRIs from the rest 57 source sites to this target domain using each method. Following this, we train a site classifier on OpenBHB training set (with $3,227$ MRIs) and validate it on the validation set (with $757$ MRIs). The site classifier is implemented based on the ResNet18 [15] architecture. Its performance on raw MRI data is recorded as the Baseline. Five metrics are used: balanced accuracy (BACC), accuracy (ACC), area under the ROC curve (AUC), F1 score (F1), and sensitivity (SEN). The results of site classification are reported in Table 1, where a lower value signifies a method’s better capability to eliminate site-related variations, thereby making it challenging for the site classifier to learn site-related features that could differentiate each site.
As indicated in Table 1, our DLEST method consistently surpasses the five competing methods by a substantial margin. For instance, after the data is harmonized by DLEST, there is a decrease in BACC by $0.340$ compared to the Baseline, and a decrease of $0.153$ when compared to the second-best StarGAN method. When compared with the three SOTA methods (i.e., CycleGAN, StarGAN, and ImUnity), our method consistently yields superior results across five metrics. This result is consistent with the qualitative result in Task 1 (see Fig. 2).
Several factors may contribute to the significant performance difference between our DLEST and the three SOTA methods. Unlike CycleGAN, which relies solely on adversarial learning between discriminators and the generators in image space to generate images resembling the target, our DLEST achieves style translation by directly manipulating the latent code of an image via EBM. This direct manipulation of latent code allows our DLEST to have fine-grained control over style translation. Furthermore, DLEST captures the underlying data distribution of the entire target site when translating an image to the target style. In contrast, StarGAN treats each target image as a unique “domain” with its own style and depends on a single reference image during style translation. Similarly, while ImUnity learns an unbiased latent representation devoid of site information, it still necessitates a single reference image for “contrast correction”. These quantitative results further suggest the superiority of DLEST in capturing the latent distribution of the target domain for effective style translation.
Figure 4: Segmentation results of CSF, GM, and WM on SRPBS in terms of (left) Dice coefficient and (right) Jaccard index. Each U-Net is trained on the target site COI and validated on the source site HUH from SRPBS.
Figure 5: Segmentation of brain tissues (i.e., CSF, GM, WM) using unharmonized MRI data (Baseline) and data harmonized by CycleGAN (top competing method) and DLEST (ours). White pixels denote the correct segmentation, red pixels denote under-segmentation, and blue pixels denote over-segmentation.
Table 1: Multi-class (58 categories) site classification results on OpenBHB, where ‘*’ denotes that the difference between a specific competing method and DLEST is statistically significant ($p<0.05$ via paired $t$-test).
[…]
Baseline $0.676$$\pm$$0.137$ $0.722$$\pm$$0.147$ $0.979$$\pm$$0.022$ $0.713$$\pm$$0.146$ $0.722$$\pm$$0.147$
HM $0.619$$\pm$$0.133$ $0.696$$\pm$$0.137$ $0.981$$\pm$$0.019$ $0.685$$\pm$$0.129$ $0.696$$\pm$$0.137$
SSIMH $0.673$$\pm$$0.083$ $0.659$$\pm$$0.076$ $0.983$$\pm$$0.008$ $0.656$$\pm$$0.079$ $0.659$$\pm$$0.076$
CycleGAN $0.588$$\pm$$0.081$ $0.660$$\pm$$0.088$ $0.974$$\pm$$0.007$ $0.645$$\pm$$0.088$ $0.660$$\pm$$0.088$
StarGAN $0.489$$\pm$$0.201$ $0.562$$\pm$$0.223$ $0.951$$\pm$$0.047$ $0.543$$\pm$$0.215$ $0.562$$\pm$$0.223$
ImUnity $0.569$$\pm$$0.074$ $0.658$$\pm$$0.066$ $0.978$$\pm$$0.008$ $0.636$$\pm$$0.066$ $0.658$$\pm$$0.066$
DLEST (Ours) $\bm{0.336}$$\pm$$0.057$ $\bm{0.483}$$\pm$$0.065$ $\bm{0.938}$$\pm$$0.020$ $\bm{0.474}$$\pm$$0.069$ $\bm{0.483}$$\pm$$0.065$
Figure 6: Synthesized MRI with site-specific styles of 3 sites from SRPBS (top 3 rows) and 1 site from OpenBHB (bottom row). The synthesized MRIs have diverse anatomical structures with consistent image styles that match the reference MRI from a specific site.
4.2.3 Task 3: Brain Tissue Segmentation
This experiment is designed to assess the effectiveness of the DLEST on a downstream brain tissue (i.e., WM, GM, and CSF) segmentation task. We begin by selecting a source site (i.e., HUH) and a target site (i.e., COI) within the SRPBS dataset and harmonize the source MRIs to the target domain using DLEST and each competing method. Similar to [23], a U-Net [32] segmentation model is trained on MRIs from the target site (COI) and directly validated on the source site (HUH). As our primary objective is to assess the impact of each harmonization method on tissue segmentation, the quality of the ground-truth segmentation is not our main concern. We thus utilize the segmentation masks obtained using the FAST pipeline in FSL as our ground truth for UNet training, providing a fair basis for comparing each method. For each harmonization method, the U-Net is trained for 100 epochs, with the mean and standard deviation of the validation Dice coefficient and the Jaccard index reported for the final 40 epochs.
Figure 4 quantitatively illustrates that despite the source and target sites sharing the same subject cohort, the direct application of a UNet model trained on the site COI to raw data from site HUH results in less than optimal segmentation. This suggests that site-related variations hinder the generalizability of the segmentation model. Our DLEST method consistently surpasses all non-learning methods (i.e., HM and SSIMH), suggesting that the latent space style translation is more effective than merely aligning the global intensity distribution. Compared with the three SOTA methods, our DLEST consistently delivers improved segmentation performance in terms of average segmentation results specifically in the segmentation of CSF and GM. While CycleGAN slightly outperforms DLEST in WM segmentation, our DLEST exhibits a significantly lower standard deviation. This indicates that the DLEST produces a more stable performance compared with CycleGAN.
In addition, Fig. 5 presents the comparative visualization of brain tissue segmentation results. The first column displays the ground truth (GT) segmentation maps, derived from FSL. The subsequent columns represent the results of three identically structured U-Nets, all trained on the target domain (i.e., COI) MRIs. In particular, the Baseline column shows validation results on unharmonized source domain (i.e., HUH) MRIs. The CycleGAN column presents the validation results on the source domain MRIs, post-harmonization to the target domain using the top competing method (i.e., CycleGAN). Similarly, the DLEST column depicts the validation results on source domain MRIs harmonized using our method. In this figure, the segmentation map uses white pixels to indicate accurate segmentation, while red and blue pixels represent under-segmentation and over-segmentation, respectively. Notably, the U-Net trained on DLEST-harmonized data demonstrated significant improvements, particularly in reducing under-segmentation of white matter near tissue boundaries and over-segmentation of gray matter near the thalamus region.
4.2.4 Task 4: MRI Synthesis
This task aims to demonstrate our framework’s ability to synthesize brain MRIs with site-specific styles. We select three sites from SRPBS and one site from OpenBHB for illustration. We evaluate the synthesized MRIs qualitatively via sample visualization and quantitatively using structural similarity index measurement (SSIM) [43], peak signal-to-noise ratio (PSNR), root mean square error (RMSE), and intensity Pearson correlation coefficient (PCC). In our analysis, we apply a binary brain mask to eliminate any influence of the background pixels when calculating each metric. As shown in Fig. 6, our DLEST is able to synthesize new brain MRIs with diverse brain shapes and anatomical structures while preserving the site-specific styles that match the given reference MRI scans. This latent space MRI synthesis is cost-efficient and helpful in training non-disease-specific learning models that require large quantities of MRI samples while the actual MRI data is limited, such as brain segmentation models.
Table 2: Quantitative result of the image-level metrics. The term “Original MRI” denotes mean and standard deviation between every pair of original images; and “Synthesized MRI” denotes mean and standard deviation between every pair of original and synthesized images.
[…]
COI Original MRI $0.900$$\pm$$0.025$ $27.206$$\pm$$2.266$ $0.045$$\pm$$0.010$ $0.916$$\pm$$0.037$
Synthesized MRI $0.901$$\pm$$0.019$ $26.563$$\pm$$1.817$ $0.048$$\pm$$0.010$ $0.925$$\pm$$0.028$
YC1 Original MRI $0.900$$\pm$$0.029$ $25.386$$\pm$$2.672$ $0.056$$\pm$$0.015$ $0.933$$\pm$$0.033$
Synthesized MRI $0.903$$\pm$$0.019$ $25.584$$\pm$$1.852$ $0.054$$\pm$$0.011$ $0.942$$\pm$$0.023$
UTO Original MRI $0.904$$\pm$$0.026$ $27.041$$\pm$$2.572$ $0.046$$\pm$$0.011$ $0.905$$\pm$$0.041$
Synthesized MRI $0.905$$\pm$$0.018$ $27.249$$\pm$$1.680$ $0.044$$\pm$$0.009$ $0.914$$\pm$$0.029$
Site 17 Original MRI $0.879$$\pm$$0.022$ $25.176$$\pm$$1.892$ $0.056$$\pm$$0.011$ $0.899$$\pm$$0.036$
Synthesized MRI $0.873$$\pm$$0.015$ $24.918$$\pm$$1.270$ $0.057$$\pm$$0.009$ $0.898$$\pm$$0.031$
For quantitative evaluation, we report the mean and the standard deviation of image-level metrics for synthesized MRIs at each site in Table 2. Results under the Original category are derived from comparisons between each original image and every other original image within the same site. The Synthesized category results are obtained by comparing each synthesized image with all original images from the same site. The negligible disparity across all four metrics between the original-original and original-synthesized image pairs implies that the synthesized images uphold the same quality as their original counterparts. Moreover, the slightly increased PCC observed in most sites indicates that the synthesized MRI slices exhibit a consistent style and fewer variations when compared to the original MRIs from the respective sites.
Figure 7: MRI quality assessment results achieved by DLEST and its variants on traveling subjects from SRPBS.
5 Discussion
5.1 Ablation Study
We investigate three downgraded variants of the proposed DLEST framework, namely (1) DLEST-PIX, which excludes pixel-wise loss in Eq. (5) during the SIG module training, (2) DLEST-CON, which omits the latent content loss in Eq. (10), and (3) DLEST-CYC, which does not incorporate the latent cycle-consistency loss in Eq. (10) when training the SST module for style translation. We train DLEST and its three variants on the SRPBS traveling subject dataset and evaluate their harmonization results using three image quality metrics: SSIM, PSNR, and PCC. As indicated in Fig. 7, while all ablations negatively impact the result, DLEST-PIX particularly yields the worst results, compared with its two counterparts. This implies that the pixel-wise, latent content and cycle-consistency losses are necessary to ensure a satisfying harmonization outcome, and especially, the pixel-wise loss plays a crucial role in preserving image content in DLEST.
Figure 8: Results of DLEST with different values of (a) $\alpha$ and (b) $\beta$ on the SRPBS traveling subject dataset.
5.2 Influence of Hyperparameters
We conduct several rounds of hyperparameter tuning to examine the influence of hyperparameters $\alpha$ and $\beta$ in Eq. (10) on the harmonization effect. This involves varying one parameter while maintaining the other constant. The results in Fig. 8 suggest that the model’s performance improves when the latent content loss has a balanced weight (i.e., $\alpha=1$). Additionally, the performance of DLEST appears to be less sensitive to variations in $\beta$ with optimal performance at $\beta=100$.
Figure 9: (a) Comparison of DLEST and DLEST+CS, and (b) results of DLEST+CS with different style/content ratios on the SRPBS traveling subject dataset.
5.3 Influence of Content Loss
The proposed DLEST uses a latent content loss in Eq. (8) to constrain the change of latent information while minimizing the $\mathcal{L}_{EBM}$ during style translation. This implicit regularization in latent space may cause slight alterations in anatomical details. To this end, we compare our method with its variant DLEST+CS that utilizes a separate style and content loss [11] to explicitly enforce content preservation while allowing the style to be translated freely. In DLEST+CS, a pre-trained VGG19 [35] is used to extract feature maps of given MRI slices. The content loss is calculated as the $l_{2}$ distance between the extracted feature maps of original MRI and the translated MRI in specific layers. The style loss is computed as the weighted average of $l_{2}$ distance between Gram matrices [11] of the target and translated MRIs in each layer. The Gram matrix is a correlation matrix between each feature map, where each entry $\mathcal{G}_{ij}$ is the inner product of vectorized feature maps $i$ and $j$. By including feature correlations of multiple layers, we obtain a multi-scale representation of the input image, which captures its texture characteristics without content information [11].
As shown in Fig. 9 (a), compared with DLEST, its variant DLEST+CS achieves slight improvement across all metrics, implying enhanced content preservation and style translation. We also investigate the influence of different ratios of style and content loss as suggested in [11]. The results in Fig. 9 (b) indicate that an emphasis on content loss (lower style/content loss ratio) leads to better harmonization performance. Despite the enhancements by DLEST+CS, it’s important to note that these come with a substantial rise in computational expense. This is primarily due to the computation of style and content loss in the ambient image space. Consequently, after each iteration of Stochastic Gradient Langevin Dynamics (SGLD), the latent code must be decoded. Furthermore, this process necessitates the use of an additional pre-trained feature extractor, adding to the complexity and computational demand.
Figure 10: Histogram of harmonized MRIs using four different target sites (i.e., YC1, UTO, HUH, and HKH) on SRPBS. Given a specific target site, MRIs from the remaining 10 sites are used as source data. The PSNR value denotes the intra-site variation of each target site.
5.4 Influence of Target Site
We further examine the impact of target site selection on harmonization using four different target sites (i.e., YC1, UTO, HUH, and HKH) in the SRPBS dataset. Similar to the main experiments, given a selected target site, we harmonize the remaining 10 sites as the source domain. Notably, only the SST module requires fine-tuning with each target change, typically requiring about 100 2D MRI slices from the target domain. This process incurs minimal computational cost, as detailed in Section 5.5. The histogram results of the harmonized MRIs, depicted in Fig. 10, underscore the influence of target site selection on the harmonization outcome. When a target site with lower intra-site variations (higher PSNR), such as site HKH, is used, the harmonized histograms across all sites align more effectively. Conversely, a target site with a lower PSNR, such as site HUH, may result in less optimal histogram alignment. This observation suggests that a target site with excessive variations could hinder the model’s ability to accurately capture the latent data distribution that characterizes the target style.
Table 3: Results of time and computation cost. For DLEST, “$a+b$” denotes the number for SIG $+$ SST. The term “Test time” is the cost of generalizing on an unseen dataset. M: Million; GMac: Giga multiply-accumulate operations; H: Hour.
Method ##\##Parameters (M) FLOPs (GMac) Training Time (H) Test Time (H) Need Re-training
CycleGAN $28.3$ $1794.1$ $29.1$ $8.0$ $Yes$
StyleGAN $161.3$ $4865.3$ $53.8$ $9.1$ $Yes$
ImUnity $671.0$ $246.2$ $25.3$ $1.5$ $No$
DLEST (Ours) $21.7$+$5.3$ $2.1$+$0.3$ $20.0$+$0.4$ $0.4$ $No$
5.5 Computational Cost Evaluation
To quantitatively assess the enhancements in computational and time efficiency, we compare our DLEST framework with three GAN-based SOTA methods across several metrics. These include the number of trainable parameters, total floating point operations (FLOPs) per forward pass, training time to convergence on the OpenBHB training dataset, testing time on the SRPBS dataset, and the necessity for retraining when applied to an unseen dataset. As demonstrated in Table 3, our model, encompassing both the SIG and SST modules, contains the fewest parameters, requires significantly fewer FLOPs per forward pass, and generally requires less training time. Furthermore, in comparison to ImUnity, which can also generalize to unseen data without retraining, our DLEST framework requires less test time to fine-tune solely the SST module. These results collectively suggest that our DLEST can achieve highly efficient and generalizable MRI harmonization.
5.6 Limitations and Future Work
While extensive experiments demonstrate that the proposed DLEST is competitive over SOTA methods, it is important to acknowledge some limitations. First, our current implementation focuses on 2D-level MRI slices, potentially overlooking volumetric variations, and can only synthesize separate MRI slices instead of the whole volume. One possible solution is to extend our implementation to 2.5D, incorporating a conditioning mechanism in the autoencoder to generate MRI slices by conditioning on adjacent slices [28]. Second, our current autoencoder employs a one-dimensional latent code to achieve the most efficient encoding and style translation. However, the aggressive 1D encoding discards some inherent spatial structures of the original image, leading to some minor loss of anatomical structures during reconstruction. It is interesting to employ 2D latent encoding [31] with a relatively mild compression rate to better preserve anatomical structures of input MRIs, which will be one of our future works. In addition, our current work only uses latent information from images, without considering biological features, such as gender, age, and relevant clinical diagnostic scores. As another further work, we plan to incorporate biological variables as a condition to improve the fidelity of generated images.
6 Conclusion
This paper proposes a novel unpaired image-level harmonization framework for multi-site MRI data. By disentangling site-invariant image generation and site-specific style translation, our model can generalize on new sites without any retraining on original data. The style translation in latent space efficiently harmonizes multiple source sites to the target site. Moreover, the model can synthesize diverse MRIs in a site-specific style without additional training. Quantitative and qualitative evaluations suggest that our DLEST outperforms several state-of-the-art harmonization methods. Additionally, our framework can be employed as a pre-processing step in downstream studies, such as mitigating site effects in multi-site MRIs and synthesizing MRIs to train general models.